### v1

# Computing Minimal Absent Words and Extended Bispecial Factors with CDAWG Space

Abstract
A string $w$ is said to be a minimal absent word (MAW) for a string $S$ if $w$ does not occur in $S$ and any proper substring of $w$ occurs in $S$. We focus on non-trivial MAWs which are of length at least 2. Finding such non-trivial MAWs for a given string is motivated for applications in bioinformatics and data compression. Fujishige et al. [TCS 2023] proposed a data structure of size $\Theta(n)$ that can output the set $\mathsf{MAW}(S)$ of all MAWs for a given string $S$ of length $n$ in $O(n+|\mathsf{MAW}(S)|)$ time, based on the directed acyclic word graph (DAWG). In this paper, we present a more space efficient data structure based on the compact DAWG (CDAWG), which can output $\mathsf{MAW}(S)$ in $O(|\mathsf{MAW}(S)|)$ time with $O(\mathsf{e}_{\min})$ space, where $\mathsf{e}_{\min}$ denotes the minimum of the sizes of the CDAWGs for $S$ and for its reversal $S^{R}$. For any strings of length $n$, it holds that $\mathsf{e}_{\min}<2n$, and for highly repetitive strings $\mathsf{e}_{\min}$ can be sublinear (up to logarithmic) in $n$. We also show that MAWs and their generalization minimal rare words have close relationships with extended bispecial factors, via the CDAWG.
1 Introduction
A string $w$ is said to be a minimal absent word (MAW) for a string $S$ if $w$ does not occur in $S$ and any proper substring of $w$ occurs in $S$. Throughout this paper, we focus on MAWs which are of form $aub$, where $a,b$ are characters and $u$ is a (possibly empty) string. Finding such MAWs for a given string is motivated for applications including bioinformatics [1, 12, 23, 19] and data compression [14, 15, 3].
Fujishige et al. [16] proposed a data structure of size $\Theta(n)$ that can output the set $\mathsf{MAW}(S)$ of all MAWs for a given string $S$ of length $n$ in $O(n+|\mathsf{MAW}(S)|)$ time, based on the directed acyclic word graph (DAWG[9]. The DAWG for string $S$, denoted $\mathsf{DAWG}(S)$, is the smallest DFA recognizing all suffixes of $S$.
In this paper, we present a more space efficient data structure based on the compact DAWG (CDAWG[10], which takes $O(\mathsf{e}_{\min})$ space and can output $\mathsf{MAW}(S)$ in $O(|\mathsf{MAW}(S)|)$ time, where $\mathsf{e}_{\min}=\{\mathsf{el}(S),\mathsf{er}(S)\}$, with $\mathsf{el}(S)$ and $\mathsf{er}(S)$ being the numbers of edges of the CDAWGs for $S$ and for its reversal $S^{R}$, respectively. For any string $S$ of length $n$ it holds that the number of edges of the CDAWG for $S$ is less than $2n$ and it can be sublinear (up to logarithmic) in $n$ for highly repetitive strings [25, 24, 5].
Our new data structure with the CDAWG is built on our deeper analysis on how the additive $O(n)$ term is required in the output time of Fujishige et al.’s DAWG-based algorithm [16]. Let $x$ be a node of $\mathsf{DAWG}(S)$ and let $u$ be the longest string represented by $x$. Suppose that there is a node $y$ of which the suffix link points to $x$, and let $a$ be the character such that $au$ is the shortest string represented by $y$. It follows that $aub$ is a MAW for $S$ if and only if node $x$ has an out-edge labeled $b$ but node $y$ does not. Fujishige et al.’s algorithm finds such characters $b$ by comparing the out-edges of $y$ and $x$ in sorted order. However, Fujishige et al.’s algorithm performs redundant comparisons even for every pair of unary nodes $x,y$ connected by a suffix link. This observation has led us to the use of the CDAWG for $S$, where unary nodes are compacted. Still, we need to manage the case where $y$ is unary but $x$ is not unary. We settle this issue by a non-trivial use of the extended longest path tree of the CDAWG [18].
In the previous work by Belazzougui and Cunial [5], they proposed another CDAWG-based data structure of size $O(\max\{\mathsf{el}(S),\mathsf{er}(S)\})$ that can output all MAWs for $S$ in $O(\max\{\mathsf{el}(S),\mathsf{er}(S)\}+|\mathsf{MAW}(S)|)$ time. The $\mathsf{el}(S)$ term comes from the Weiner links for all nodes of the CDAWG, which can be as large as $\Omega(\mathsf{er}(S)\sqrt{n})$ for some strings [18]. Note that our data structures uses only $O(\mathsf{e}_{\min})=O(\min\{\mathsf{el}(S),\mathsf{er}(S)\})$ space and takes only $O(|\mathsf{MAW}(S)|)$ time to report the output. We also present a simpler algorithm that restores the string $u$ in linear time in its length using the CDAWG-grammar [5], without using level-ancestor data structures [8, 7], which can be of independent interest.
We also show how our CDAWG-based data structure of $O(\mathsf{e}_{\min})$ space can output all extended bispecial factors (EBFs) [2] for the input string $S$ in output optimal time. This proposed method can be seen as a more space-efficient variant of Almirantis et al.’s method [2] that is built on suffix trees [26]. We also show that minimal rare words (MRWs[4], which are generalizations of MAWs and minimal unique substrings MUSs [17], have close relationships with EBFs through the CDAWG.
2 Preliminaries
2.1 Basic Notations
Let $\Sigma$ be an alphabet. An element of $\Sigma$ is called a character. An element of $\Sigma^{\ast}$ is called a string. The length of a string $S$ is denoted by $|S|$. The empty string $\varepsilon$ is the string of length 0. Let $\Sigma^{+}=\Sigma^{*}\setminus\{\varepsilon\}$. If $S=xyz$, then $x$, $y$, and $z$ are called a prefix, substring (or factor), and suffix of $S$, respectively. Let $\mathsf{Prefix}(S)$, $\mathsf{Substr}(S)$, and $\mathsf{Suffix}(S)$ denote the sets of the prefixes, substrings, and suffixes of $S$, respectively. An element in $\mathsf{Suffix}(S)\setminus\{S\}$ is called a proper suffix of $S$. For any $1\leq i\leq|S|$, $S[i]$ denotes the $i$-th character of string $S$. For any $1\leq i\leq j\leq|S|$, $S[i..j]$ denotes the substring of $S$ that begins at position $i$ and ends at position $j$. For convenience, let $T[i..j]=\varepsilon$ for any $i>j$. We say that string $w$ occurs in string $S$ iff $w$ is a substring of $S$. Let $\mathsf{occ}_{S}(w)=|\{i\mid w=S[i..i+|w|-1]\}|$ denote the number of occurrences of $w$ in $S$. For convenience, let $\mathsf{occ}_{S}(\varepsilon)=|S|+1$. For a string $S$, let $S^{R}=S[|S|]\cdots S[1]$ denote the reversed string of $S$.
2.2 MAW, MUS, MRW, and EBF
A string $w$ is said to be an absent word for another string $S$ if $\mathsf{occ}_{S}(w)=0$. An absent word $w$ for $S$ is called a minimal absent word (MAW) for $S$ if $\mathsf{occ}_{S}(w[1..|w|-1])\geq 1$ and $\mathsf{occ}_{S}(w[2..|w|])\geq 1$. By definition, any character $c$ not occurring in $S$ is a MAW for $S$. In the rest of this paper, we exclude such trivial MAWs and we focus on non-trivial MAWs of form $aub$, where $a,b\in\Sigma$ and $u\in\Sigma^{*}$. We denote by $\mathsf{MAW}(S)$ the set of all non-trivial MAWs for $S$.
[…]
A string $w$ is called a minimal rare word MRW for another string $S$ if
  • (1)
    $w=c\in\Sigma$, or
  • (2)
    $w=aub$ with $a,b\in\Sigma$ and $u\in\Sigma^{*}$, such that $\mathsf{occ}_{S}(au)>\mathsf{occ}_{S}(aub)$ and $\mathsf{occ}_{S}(ub)>\mathsf{occ}_{S}(aub)$.
In particular, an MRW $w$ for $S$ is a MAW for $S$ when $\mathsf{occ}_{S}(w)=0$. Also, an MRW $w$ is a MUS for $S$ when $\mathsf{occ}_{S}(w)=1$. Let $\mathsf{MRW}(S)$ denote the set of non-trivial MRWs of type (2) for string $S$. Then, $\mathsf{MAW}(S)\subseteq\mathsf{MRW}(S)$ and $\mathsf{MUS}(S)\subseteq\mathsf{MRW}(S)$ hold.
A string $w=aub$ with $a,b\in\Sigma$ and $u\in\Sigma^{*}$ is called an extended bispecial factor (EBF) for another string $S$ if $\mathsf{occ}_{S}(a^{\prime}u)\geq 1$ and $\mathsf{occ}_{S}(ub^{\prime})\geq 1$ for some characters $a^{\prime}\neq a$ and $b^{\prime}\neq b$. Let $\mathsf{EBF}(S)$ denote the set of EBFs for string $S$.
2.3 Maximal Substrings and Maximal Repeats
A substring $u$ of a string $S$ is said to be left-maximal in $S$ if (1) there are distinct characters $a,b$ such that $au,bu\in\mathsf{Substr}(S)$, or (2) $u\in\mathsf{Prefix}(S)$. Symmetrically, a substring $u$ of $S$ is said to be right-maximal in $S$ iff (1) there are distinct characters $a,b$ such that $ua,ub\in\mathsf{Substr}(S)$, or (2) $u\in\mathsf{Suffix}(S)$. A substring $u$ of $S$ is said to be maximal in $S$ iff $u$ is both left-maximal and right-maximal in $S$. A maximal substring $u$ is said to be a maximal repeat if $\mathsf{occ}_{S}(u)\geq 2$. Let $\mathsf{MRep}(S)$ denote the set of maximal repeats in $S$. For a maximal repeat $u$ of a string $S$, the characters $a\in\Sigma$ such that $\mathsf{occ}_{S}(au)\geq 1$ are called the left-extensions of $u$, and the characters $b\in\Sigma$ such that $\mathsf{occ}_{S}(ub)\geq 1$ are called the right-extensions of $u$.
The following relationship between MAWs and maximal repeats is known:
Lemma 1 ([4] and Theorem 1 of [22]).
For any MRW $aub$ for string $S$ with $a,b\in\Sigma$ and $u\in\Sigma^{*}$, $u$ is a maximal repeat in $S$.
2.4 CDAWG
Figure 1: $\mathsf{CDAWG}(S)$ for string $S=\mathtt{ababcbababcbc\$}$.
The compact directed acyclic word graph (CDAWG) of string $S$, that is denoted $\mathsf{CDAWG}(S)=(\mathsf{V},\mathsf{E})$, is a path-compressed smallest partial DFA that represents $\mathsf{Suffix}(S)$, such that there is a one-to-one correspondence between the nodes of $\mathsf{CDAWG}(S)$ and the maximal substrings in $S$. Intuitively, $\mathsf{CDAWG}(S)$ can be obtained by merging isomorphic subtrees of the suffix tree [26] for $S$. Therefore, each string represented by a node $u$ of $\mathsf{CDAWG}(S)$ is a suffix of all the other longer strings represented by the same node $u$. For convenience, for any CDAWG node $u\in\mathsf{V}$, let $\mathsf{str}(u)$ denote the longest string represented by the node $u$. The suffix link of a CDAWG node $u$ points to another CDAWG node $v$ iff $\mathsf{str}(v)$ is the longest proper suffix of $\mathsf{str}(u)$ that is not represented by $u$. We denote it by $\mathsf{slink}(u)=v$. See Fig. 1 for an example of $\mathsf{CDAWG}(S)$.
We remark that every internal node in $\mathsf{CDAWG}(S)$ corresponds to a maximal repeat in $S$, that is, $\mathsf{MRep}(S)=\{\mathsf{str}(v)\mid\mbox{$v$ is an internal node of $\mathsf{CDAWG}(S)$}\}$. The first characters of the labels of the out-going edges of an internal node of $\mathsf{CDAWG}(S)$ are the right-extensions of the corresponding maximal substring for that node. Hence the number of edges in $\mathsf{CDAWG}(S)$ is equal to the total number of right-extensions of maximal repeats in $\mathsf{MRep}(S)$, which is denoted by $\mathsf{er}(S)$. Similarly, we denote by $\mathsf{el}(S)$ the number of left-extensions of the maximal repeats in $\mathsf{MRep}(S)$. Blumer et al. [10] showed that $\min\{\mathsf{er}(S),\mathsf{el}(S)\}<2n$ for any string $S$ of length $n$. On the other hand, it is known that there is an $\Omega(\sqrt{n})$ gap between $\mathsf{er}(S)$ and $\mathsf{el}(S)$:
Lemma 2 (Lemma 9 of [18]).
There exists a family of strings $S$ of length $n$ such that $\mathsf{er}(S)=\Theta(\sqrt{n})$ and $\mathsf{el}(S)=\Theta(n)$.
The length of an edge $(u,y,v)\in\mathsf{V}\times\Sigma^{+}\times\mathsf{V}$ in $\mathsf{CDAWG}(S)$ is its string label length $|y|$. An edge $(u,y,v)$ is called a primary edge if $|\mathsf{str}(u)|+|y|=|\mathsf{str}(v)|$, and $(u,y,v)$ is called a secondary edge otherwise ($|\mathsf{str}(u)|+|y|<|\mathsf{str}(v)|$). Namely, the primary edges are the edges in the longest paths from the source to all the nodes of $\mathsf{CDAWG}(S)$. Thus, there are exactly $|\mathsf{V}|-1$ primary edges. Let $\mathsf{LPT}(S)$ denote the spanning tree of $\mathsf{CDAWG}(S)$ that consists of the primary edges, i.e., $\mathsf{LPT}(S)$ is the tree of the longest paths from the source to all nodes of $\mathsf{CDAWG}(S)$.
Since $\mathsf{MRep}(S^{R})=\{w^{R}\mid w\in\mathsf{MRep}(S)\}$ holds for any string $S$, the CDAWGs for $S$ and $S^{R}$ can share the same nodes. Let $\mathsf{E}^{R}$ denote the set of edges of $\mathsf{CDAWG}(S^{R})$. The primary (resp. secondary) edges in $\mathsf{E}^{R}$ are also called hard (resp. soft) Weiner links of $\mathsf{CDAWG}(S)$. Each Weiner link is labeled by the first character of the label of the corresponding edge in $\mathsf{E}^{R}$. Note that $|\mathsf{E}^{R}|=\mathsf{el}(S)$.
3 CDAWG Grammars
A grammar-compression is a method that represents an input string $S$ with a context-free grammar $\mathcal{G}$ that only generates $S$. The size of $\mathcal{G}$ is evaluated by the total length of the right-hand sides of the productions, which is equal to the size of the minimal DAG obtained by merging isomorphic subtrees of the derivation tree for $\mathcal{G}$. Hence the more repeats $S$ has, the smaller grammar $S$ tends to have.
Assume that $S$ terminates with a unique character \$currency-dollar\\$\$ not occurring elsewhere in $S$. Belazzougui and Cunial [5] proposed a grammar-compression based on the CDAWGs. We use a slightly modified version of their grammar: The CDAWG-grammar for string $S$ of length $n$, denoted $\mathcal{G}_{\mathsf{CDAWG}}$, is such that:
  • The non-terminals are the nodes in $\mathsf{CDAWG}(S)$;
  • The root of the derivation tree for $\mathcal{G}_{\mathsf{CDAWG}}$ is the sink of $\mathsf{CDAWG}(S)$;
  • The derivation tree path from its root to its leaf representing the $i$-th character $S[i]$ is equal to the reversed path from the sink to the source of $\mathsf{CDAWG}(S)$ that spells out the $i$-th (text positioned) suffix $S[i..n]$.
By definition, the DAG for the CDAWG-grammar $\mathcal{G}_{\mathsf{CDAWG}}$ is isomorphic to the reversed DAG for $\mathsf{CDAWG}(S)$, and hence, the size for $\mathcal{G}_{\mathsf{CDAWG}}$ is $\mathsf{er}(S)$. See Fig. 2 for an example.
Belazzougui and Cunial [5] showed how to extract the string label $x$ of a given CDAWG edge in $O(|x|)$ time in a sequential manner (i.e. from $x[1]$ to $x[|x|]$), by augmenting the CDAWG-grammar with a level-ancestor data structure [8, 7] that works on the word RAM. This allows one to represent $\mathsf{CDAWG}(S)$ in $O(\mathsf{er}(S))$ space and perform pattern matching queries in optimal time, without explicitly storing string $S$. They also used this technique to output a MAW $aub$ as a string: Since $u$ is a maximal repeat, it suffices to decompress the path labels in $\mathsf{LPT}(S)$ from the source to the node $v$ that corresponds to $u$ (i.e. $\mathsf{str}(v)=u$).
We show that, when our goal is simply to obtain the longest string $\mathsf{str}(v)$ for a given CDAWG node $v$, then level-ancestor data structures are not needed.
Figure 2: $\mathsf{CDAWG}(S)$ (left) and the CDAWG-grammar $\mathcal{G}_{\mathsf{CDAWG}}$ (right) for the running example from Fig. 1. The dashed arcs represent the suffix links of the nodes of $\mathsf{CDAWG}(S)$. To obtain $\mathsf{str}(X_{3})=\mathtt{ababcb}$ for the CDAWG node $X_{3}$, we first decompress the non-terminal $X_{3}$ and obtain $\mathtt{ababc}$. We move to the node $X_{2}$ by following the suffix link of $X_{3}$. We then decompress the non-terminal $X_{2}$ and obtain the remaining $\mathtt{b}$.
Lemma 3.
Given $\mathsf{CDAWG}(S)=(\mathsf{V},\mathsf{E})$ with suffix links and grammar $\mathcal{G}_{\mathsf{CDAWG}}$ of total size $O(\mathsf{er}(T))$, one can decompress the longest string $\mathsf{str}(v)$ represented by a given node $v$ in $O(|\mathsf{str}(v)|)$ time.
Proof.
For each CDAWG node $v\in\mathsf{V}$, let $X(v)$ denote the non-terminal in the grammar $\mathcal{G}_{\mathsf{CDAWG}}$ that corresponds to $v$. We first decompress $X(v)$ and let $p_{v}$ be the decompressed string. Then, the length of $p_{v}$ is equal to the number of (reversed) paths between the source and $v$ in the CDAWG. Thus, if the suffix link $\mathsf{slink}(v)$ points to the source, then $p_{v}$ is the whole string $\mathsf{str}(v)$. Otherwise, then let $v_{0}=v$. We take the suffix link $\mathsf{slink}(v_{0})=v_{1}$, obtain string $p_{v_{1}}$ by decompressing $X(v_{1})$. This gives us a longer prefix $p_{v_{0}}p_{v_{1}}$ of $\mathsf{str}(v)$, since in any interval $[i..i+|\mathsf{str}(v)|-1]$ of leaves of the derivation tree for $\mathcal{G}_{\mathsf{CDAWG}}$ that represents $\mathsf{str}(v)$, $p_{v_{1}}$ immediately follows $p_{v_{0}}$. We continue this process until we encounter the last node $v_{k}$ in the suffix link chain from $v$, of which the suffix link points to the source. Then, the concatenated string $p_{v_{0}}p_{v_{1}}\cdots p_{v_{k}}$ is $\mathsf{str}(v)$.
When there are no unary productions of form $X\rightarrow Y$, then the total number of nodes visited for obtaining $p_{v_{0}}p_{v_{1}}\cdots p_{v_{k}}=\mathsf{str}(v)$ is linear in its length. Otherwise, then for each maximal sequence of unary productions $X_{1}\rightarrow X_{2}\rightarrow\cdots\rightarrow X_{\ell}\rightarrow Y$, we create a pointer from $X_{i}$ to $Y$ for every $1\leq i\leq\ell$ so that we can extract $\mathsf{str}(v)$ in time linear in its length [^1] A similar technique was used in the original CDAWG-grammar by Belazzougui and Cunial [5], where they explicitly removed such unary productions.. ∎
4 Computing MAWs and EBFs with CDAWG
4.1 Computing MAWs
In this section, we first show the following:
Theorem 1.
There exists a data structure of $O(\mathsf{e}_{\min})$ space that can output all MAWs for string $S$ in $O(|\mathsf{MAW}(S)|)$ time with $O(1)$ working space, where $\mathsf{e}_{\min}=\min\{\mathsf{er}(S),\mathsf{el}(S)\}$.
For ease of description of our algorithm, let us assume that our input string $S$ begins and ends with unique terminal symbols $\sharp$ and \$currency-dollar\\$\$ not occurring inside $S$, respectively. We also assume ,\$$\sharp,\$\in\Sigma$ so that they can be the first or the last characters for MAWs and EBFs for $S$, respectively. The case without $\sharp$ and \$currency-dollar\\$\$ can be treated similarly.
We remark that MAWs have symmetric structures such that $aub\in\mathsf{MAW}(S)$ iff $bu^{R}a\in\mathsf{MAW}(S^{R})$. This implies that one can work with the smaller one of the two: $\mathsf{CDAWG}(S)$ of $O(\mathsf{er}(S))$ space and $\mathsf{CDAWG}(S^{R})$ of $O(\mathsf{el}(S))$ space. In what follows, we assume w.l.o.g. that $\mathsf{e}_{\min}=\mathsf{er}(S)$, unless otherwise stated.
On top of the edge-sorted $\mathsf{CDAWG}(S)$ augmented with suffix links and the CDAWG-grammar $\mathcal{G}_{\mathsf{CDAWG}}$ for $S$, we use the extended version of $\mathsf{LPT}(S)$ proposed in [18], that is denoted $\mathsf{LPT}^{+}(S)$. Let $v$ be any node in the longest path tree $\mathsf{LPT}(S)$, and suppose that its corresponding node in $\mathsf{CDAWG}(S)$ has a secondary out-going edge $(v,u)$. Then, we create a copy $\ell$ of $u$ as a leaf, and add a new edge $(v,\ell)$ to the tree. $\mathsf{LPT}^{+}(S)$ is the tree obtained by adding all secondary edges of $\mathsf{CDAWG}(S)$ to $\mathsf{LPT}(S)$ in this way, and thus, $\mathsf{LPT}^{+}(S)$ has exactly $\mathsf{er}(S)$ edges. For any node $v$ in $\mathsf{LPT}^{+}(S)$, let $\mathsf{str}(v)$ denote the string label in the path from the root to $v$.
Let $u$ be any node in $\mathsf{CDAWG}(S)$ and suppose it has $k$ in-coming edges. The above process of adding secondary edges to $\mathsf{LPT}(S)$ can be regarded as splitting $u$ into $k+1$ nodes $u^{\prime},\ell_{1},\ldots,\ell_{k}$, such that $|\mathsf{str}(u^{\prime})|<|\mathsf{str}(\ell_{1})|<\cdots<|\mathsf{str}(\ell_{k})|$. We define the suffix links of these nodes such that $\mathsf{slink}(u^{\prime})=\ell_{1}$, $\mathsf{slink}(\ell_{i-1})=\ell_{i}$ for $1\leq i<k$, and $\mathsf{slink}(\ell_{k})=v^{\prime}$ such that $v^{\prime}$ is the node in $\mathsf{LPT}^{+}(S)$ that corresponds to the destination node $v=\mathsf{slink}(u)$ of the suffix link of $u$ in $\mathsf{CDAWG}(S)$. We call these copied nodes $\ell_{1},\ldots,\ell_{k}$ as gray nodes, and the other node $u^{\prime}$ as a white node. For convenience, we assume that the source is a white node. See Fig. 3 for illustration.
Figure 3: $\mathsf{LPT}^{+}(S)$ for the running example from Fig. 1 and 2. The double-lined arcs represent primary edges, and the single-lined arcs represent secondary edges. For edge $(\hat{v},\hat{u})$ with string label $\mathtt{c\$}$, consider the path $\langle v,u_{1}\rangle$ that spells out $\mathtt{c\$}$ and is obtained by the fast link. The CDAWG node that corresponds to $u_{2}=\mathtt{bc}$ has a virtual soft Weiner link with label $\mathtt{c}$ pointing to the CDAWG node that corresponds to $\hat{u}$. Node $u_{2}$ has an out-edge with $\mathtt{b}$. Therefore, $\mathtt{c}u_{2}\mathtt{b}=\mathtt{cbcb}$ is a MAW for the string $S=\mathtt{ababcbababcbc\$}$.
In what follows, we show how to compute MAWs for a given edge $(\hat{v},\hat{u})$ in $\mathsf{LPT}^{+}(S)$. Note that $\hat{v}$ is always a white node since every gray node has no children. Let $X$ be the string label of $(\hat{v},\hat{u})$. Let $v$ be the first white node in the chain of suffix links from $\hat{v}$ and let $u$ be the descendant of $v$ in $\mathsf{LPT}^{+}(S)$ such that the path from $v$ to $u$ spells out $X$. We maintain a fast link from the edge $(\hat{v},\hat{u})$ to the path $\langle v,u\rangle$ so we can access from $(\hat{v},\hat{u})$ to the terminal nodes $v$ and $u$ in the path in $O(1)$ time. We have the four following cases (See also Fig. 4):
  1. (A)
    $\hat{u}$ and $u$ are both white nodes;
  2. (B)
    $\hat{u}$ is a white node and $u$ is a gray node;
  3. (C)
    $\hat{u}$ is a gray node and $u$ is a white node;
  4. (D)
    $\hat{u}$ and $u$ are both gray nodes.
Figure 4: Cases (A), (B), (C), and (D) for computing MAWs from a given edge $(\hat{v},\hat{u})$ on $\mathsf{LPT}^{+}(S)$. The bold arc represents the fast link from edge $(\hat{v},\hat{u})$ to path $\langle v,u\rangle$, where $u=u_{1}$. The dashed arcs represent suffix links. The dotted arcs in Cases (C) and (D) are additional pointers for $O(1)$-time access from $\hat{u}$ to $\tilde{u}$.
How to deal with Case (A): In Case (A), $u=\mathsf{slink}(\hat{u})$.
Let $m$ be the number of nodes in the path from $v$ to $u$, with $v$ exclusive. In the sequel, we consider the case where $m>1$, as the case with $m=1$ is simpler to show. Let $u_{1}~{}(=u),u_{2},\ldots,u_{m}$ denote these $m$ nodes in the path arranged in decreasing order of their depths. We process each $u_{i}$ in increasing order of $i$. For simplicity, we will identify each node $u_{i}$ as the string $\mathsf{str}(u_{i})$ represented by $u_{i}$.
For $i=1$, node $u_{1}~{}(=u)$ has a hard Weiner link to $\hat{u}$. Let $a$ be the character label of this hard Weiner link. Let $\mathsf{child}(\hat{u})$ and $\mathsf{child}(\mathsf{slink}(u))=\mathsf{child}(u_{1})$ be the sorted lists of the first characters of the out-edge labels from $\hat{u}$ and from $u_{1}$. Note that $\mathsf{child}(\hat{u})\subseteq\mathsf{child}(u_{1})$. Then, it follows from the definition of MAWs and Lemma 2 that $au_{1}b\in\mathsf{MAW}(S)$ iff $b\in\mathsf{child}(u_{1})\setminus\mathsf{child}(\hat{u})$.
For each $i=2,\ldots,m$, observe that node $u_{i}$ has a (virtual) soft Weiner link with the same character label $a$, and leading to the same node $\hat{u}$. We emphasize that this soft Weiner link only virtual and is not stored in our $O(\mathsf{er}(S))$-space data structure. We have the following lemma:
Lemma 4.
For each internal node $u_{i}$ with $2\leq i\leq m$ in the path between $v$ and $u$, there exists at least one MAW $au_{i}b$ for $S$, where $a,b\in\Sigma$.
Proof.
Let $c$ be the first character of the label of the edge $(u_{i+1},u_{i})$. Let $\hat{u}_{i}$ denote the (virtual) implicit node on the edge $(v,u)$, such that $\mathsf{slink}(\hat{u}_{i})=u_{i}$. Note that $c$ is the unique character that immediately follows the locus of $\hat{u}_{i}$ in the string label of edge $(v,u)$. Since $u_{i}$ is branching but $\hat{u}_{i}$ is non-branching, the set $\mathsf{child}(u_{i})\setminus\mathsf{child}(\hat{u}_{i})=\mathsf{child}(u_{i})\setminus\{c\}$ is not empty. Thus, for any character $b\in\mathsf{child}(u_{i})\setminus\{c\}$, we have $au_{i}b\in\mathsf{MAW}(S)$. ∎
Due to Lemma 4, we output MAWs $au_{i}b$ for every character $b\in\mathsf{child}(u_{i})\setminus\{c\}$.
How to deal with Case (B): In Case (B), $u$ is a gray node, and there might be other node(s) in the suffix link chain from $\hat{u}$ to $u$. No MAW is reported for the gray node $u_{1}$ ($=u$). The rest is the same as Case (A).
How to deal with Cases (C) and (D): In Cases (C) and (D) where node $\hat{u}$ is a gray node, we use the white node $\tilde{u}$, which corresponds to the CDAWG node from which $\hat{u}$ is copied. We output MAW(s) by comparing out-edges of $\tilde{u}$ and $u_{1}$. We store a pointer from $\hat{u}$ to $\tilde{u}$, so that we can access $\tilde{u}$ from $\hat{u}$ in $O(1)$ time (see also Fig. 4). The rest is the same as Cases (B) and (A).
4.1.1 Putting all together.
By performing the above procedure for all edges $(v,u)$ in $\mathsf{LPT}^{+}(S)$, we can output all elements in $\mathsf{MAW}(S)$.
Let us analyze the time complexity and the data structure space usage. In the case with $i\geq 2$, by Lemma 4, accessing each node $u_{i}$ can be charged to any one of the MAWs reported with $u_{i}$. The time required for $u_{i}$ is thus linear in the number of reported MAWs $au_{i}b$. In the case with $i=1$, we compare the sorted lists $\mathsf{child}(\hat{u})$ and $\mathsf{child}(u_{1})$. Each successful character comparison with a character $b\in\mathsf{child}(u_{1})\setminus\mathsf{child}(\hat{u})$ for which MAW $au_{1}b$ is reported. On the other hand, each unsuccessful character comparison with a character $b\in\mathsf{child}(u_{1})\cap\mathsf{child}(\hat{u})$ can be charged to the corresponding out-edge of $\hat{u}$ whose string label begins with $b$. Thus, we can output all MAWs in $O(\mathsf{er}(S)+|\mathsf{MAW}(S)|)$ time. Further, when $|\mathsf{MAW}(S)|\leq\mathsf{er}(S)$, then we can store the set $\mathsf{MAW}(S)$ with $O(|\mathsf{MAW}(S)|)\subseteq O(\mathsf{er}(S))$ space and output each MAW in $O(1)$ time, using Lemma 3. This leads to our $O(|\mathsf{MAW}(S)|)$-time $O(\mathsf{er}(S))$-space data structure.
To analyze the working memory usage, let us consider the number of pointers. For each given edge $(\hat{v},\hat{u})$, we climb up the path from $\mathsf{slink}(\hat{u})=u_{1}$ to $u_{m}$ one by one, until reaching $\mathsf{slink}(\hat{v})=v$, which requires $O(1)$ pointers. We also use $O(1)$ additional pointers for the character comparisons on the sorted lists of each pair of nodes. Thus the total working space is $O(1)$. We have proven Theorem 1.
4.1.2 Combinatorics on MAWs with CDAWG.
Theorem 1 also leads to the following combinatorial property of MAWs:
Theorem 2.
For any string $S$, $|\mathsf{MAW}(S)|=O(\sigma\mathsf{e}_{\min})$. This bound is tight.
Proof.
For each node $u_{i}$ of $\mathsf{LPT}^{+}(S)$, the number of distinct characters $a$ such that $au_{i}b\in\mathsf{MAW}(S)$ is at most the number of (hard or sort) Weiner links from $u_{i}$, which is bounded by $\sigma$. The number of distinct characters $b$ such that $au_{i}b\in\mathsf{MAW}(S)$ does not exceed the number of out-edges from $u_{i}$. By summing them up for all nodes $u_{i}$ of $\mathsf{LPT}^{+}(S)$, we obtain $|\mathsf{MAW}(S)|=O(\sigma\cdot\mathsf{er}(S))$. Since the same argument holds for the reversed string $S^{R}$ and $\mathsf{LPT}^{+}(S^{R})$, we have $|\mathsf{MAW}(S)|=O(\sigma\cdot\min\{\mathsf{er}(S),\mathsf{el}(S)\})=O(\sigma\mathsf{e}_{\min})$.
It is known [20] that for the de Bruijn sequence $B_{k,\sigma}$ of order $k$ over an alphabet of size $\sigma$, $\mathsf{MAW}(B_{k,\sigma})=\Omega(\sigma n)$, where $n=|B_{k,\sigma}|$. Since $\mathsf{e}_{\min}\leq\mathsf{er}(S)<2n$ for any string $S$ of length $n$, we have $\Omega(\sigma\mathsf{e}_{\min})$ for $B_{k,\sigma}$. ∎
4.2 Computing EBFs
We first prove the following combinatorial property for EBFs:
Lemma 5.
For any string $S$, $|\mathsf{EBF}(S)|\leq\mathsf{er}(S)+\mathsf{el}(S)-|\mathsf{V}|+1$.
Proof.
Recall our algorithm for computing $\mathsf{MAW}(S)$ in Section 4.1. Let $u_{i}$ be a node that has (soft or hard) Weiner link with label $a\in\Sigma$ and has an out-edge starting with $b\in\Sigma$. Assuming $S[1]=\sharp$ and $S[|S|]=\$$, there are characters $a^{\prime},b^{\prime}\in\Sigma$ such that $a^{\prime}\neq a$, $b^{\prime}\neq b$, and $a^{\prime}u_{i},u_{i}b^{\prime}\in\mathsf{Substr}(S)$. Note also that $au_{i},u_{i}b\in\mathsf{Substr}(S)$. Thus, if $au_{i}b\in\mathsf{Substr}(S)$, then $au_{i}b\in\mathsf{EBF}(S)$. If the locus for $au_{i}$ is the node $\hat{u}$, in which case the Weiner link from $u_{i}~{}(=u_{1})$ to $\hat{u}$ is hard, then the EBF $au_{i}b$ is charged to the out-edge of $\hat{u}$ that starts with $b$. If the locus for $au_{i}$ is on the edge $(\hat{v},\hat{u})$, then the EBF $au_{i}b$ is charged to the soft-Weiner link from $u_{i}$ that labeled $a$. Since the number of hard Weiner links is $|V|-1$, the number of soft-Weiner links is $\mathsf{el}(S)-|V|+1$. Thus we have $|\mathsf{EBF}(S)|=\mathsf{er}(S)+\mathsf{el}(S)-|V|+1$ for any string $S$ with $S[1]=\sharp$ and $S[|S|]=\$$.
If $S[1]\neq\sharp$ and/or $S[|S|]\neq\$$, then some node(s) $u_{i}$ in $\mathsf{CDAWG}(S)$ may only have unique characters $a,b$ such that $au_{i},u_{i}b\in\mathsf{Substr}(S)$. Thus the bound is $|\mathsf{EBF}(S)|\leq\mathsf{er}(S)+\mathsf{el}(S)-|\mathsf{V}|+1$ for arbitrary strings $S$. ∎
By modifying our algorithm for computing MAWs for $S$ as is described in the poof for Lemma 5, we can obtain the following:
Theorem 3.
There exists a data structure of $O(\mathsf{e}_{\min})$ space that can output all EBFs for string $S$ in $O(|\mathsf{EBF}(S)|)$ time with $O(1)$ working space.
4.3 Computing length-bounded MAWs and EBFs
For any set $A\subseteq\Sigma^{*}$ of strings and an integer $\ell>1$, let $A^{\geq\ell}=\{w\in A\mid|w|\geq\ell\}$ and $A^{\leq\ell}=\{w\in A\mid|w|\leq\ell\}$.
Theorem 4.
There exists a data structure of $O(\mathsf{e}_{\min})$ space which, given a query length $\ell$, can output
  • $\mathsf{MAW}(S)^{\geq\ell}$ in $O(|\mathsf{MAW}(S)^{\geq\ell}|+1)$ time,
  • $\mathsf{MAW}(S)^{\leq\ell}$ in $O(|\mathsf{MAW}(S)^{\leq\ell}|+1)$ time,
  • $\mathsf{EBF}(S)^{\geq\ell}$ in $O(|\mathsf{EBF}(S)^{\geq\ell}|+1)$ time,
  • $\mathsf{EBF}(S)^{\leq\ell}$ in $O(|\mathsf{EBF}(S)^{\leq\ell}|+1)$ time,
with $O(1)$ working space.
Proof.
We only describe how to compute length-bounded MAWs. We can compute length-bounded EBFs similarly.
In our data structure for reporting $\mathsf{MAW}(S)^{\leq\ell}$, we precompute the largest integer $m$ with $|\mathsf{MAW}(S)^{\leq m}|\leq\mathsf{e}_{\min}$, and store the elements in $\mathsf{MAW}(S)^{\leq m}$ using Lemma 3, sorted in their lengths. Given a query length $\ell$, if $\ell\leq m$, then we can output the elements of $\mathsf{MAW}(S)^{\leq\ell}$ by reporting the stored MAWs in increasing order, in $O(|\mathsf{MAW}(S)^{\leq\ell}|+1)$ time. Otherwise (if $\ell>m$), then we process each edge $(\hat{v},\hat{u})$ in $\mathsf{LPT}^{+}(S)$ in increasing order of $|\mathsf{str}(\hat{u})|$. We need to access the internal nodes in the path from $\mathsf{slink}(\hat{v})$ to $\mathsf{slink}(\hat{u})$ topdown, in $O(1)$ time each. This can be done in $O(1)$ time per node by augmenting $\mathsf{LPT}^{+}(S)$ with the level-ancestor (LA) data structure [8, 6]. We stop processing the edge $(\hat{v},\hat{u})$ as soon as we find the shallowest ancestor $u_{i}$ of $u=\mathsf{slink}(\hat{u})$ such that $|\mathsf{str}(u_{i})|>\ell-2$. Since $|\mathsf{MAW}(S)^{\leq\ell}|>\mathsf{e}_{\min}$ for any $\ell<m$, this algorithm outputs $\mathsf{MAW}(S)^{\leq\ell}$ in $O(|\mathsf{MAW}(S)^{\leq\ell}|+1)$ time.
Computing $\mathsf{MAW}(S)^{\geq\ell}$ is symmetric, except that the LA data structure is not needed: We can check the nodes between $\mathsf{slink}(\hat{u})$ and $\mathsf{slink}(\hat{v})$ in deceasing order of their lengths, in $O(1)$ time each, by simply climbing up the path. ∎
Theorem 4 is an improvement and a generalization over the previous work [11] that gave an $O(\ell n)$-space data structure that reports $\mathsf{MAW}(S)^{\leq\ell}$ in $O(\sigma n)$ time.
5 Computing MRWs
Let $\mathsf{MRW}_{k}(S)=\{w\in\mathsf{MRW}(S)\mid\mathsf{occ}_{S}(w)=k\}$. Then $\mathsf{MRW}_{0}(S)=\mathsf{MAW}(S)$ and $\bigcup_{k\geq 0}\mathsf{MRW}_{k}(S)=\mathsf{MRW}(S)$ hold. We show the following lemma. Recall that any string in $\mathsf{MAW}(S)$ or $\mathsf{EBF}(S)$ is of length at least $2$.
Lemma 6.
For any string $S$ with $S[1]=\sharp$ and $S[|S|]=\$$, $\mathsf{MRW}(S)\setminus\mathsf{EBF}(S)=\mathsf{MAW}(S)$. In other words, $\mathsf{MRW}(S)\cap\mathsf{EBF}(S)=\bigcup_{k\geq 1}\mathsf{MRW}_{k}(S)$.
Proof.
Clearly, $\mathsf{MAW}(S)\cap\mathsf{EBF}(S)=\emptyset$. It suffices to show that $\bigcup_{k\geq 1}\mathsf{MRW}_{k}(S)\subseteq\mathsf{EBF}(S)$. Let $aub$ be an element in $\mathsf{MRW}_{k}(S)$ for some $k\geq 1$ where $a,b\in\Sigma$ and $u\in\Sigma^{\star}$. Since $aub$ is a minimal rare word occurring $k$ times in $S$, $\mathsf{occ}_{S}(au)\geq k+1$ and $\mathsf{occ}_{S}(ub)\geq k+1$. Also, since $k+1\geq 2$, the last character of $au$ is not equal to \$currency-dollar\\$\$, which is a unique character. Similarly, the first character of $ub$ is not equal to $\sharp$. Thus, any occurrence of $au$ (resp., $ub$) in $S$ is not a suffix (resp., prefix) of $S$. Hence, there must be characters $b^{\prime}\neq b$ and $a^{\prime}\neq a$ such that $aub^{\prime},a^{\prime}ub\in\mathsf{Substr}(S)$. Therefore, $aub$ is an extended bispecial factor. ∎
Next, we estimate the number of minimal rare words occurring in $S$.
Lemma 7.
$|\bigcup_{k\geq 1}\mathsf{MRW}_{k}(S)|\leq\mathsf{e}_{\min}$.
Proof.
For any $w\in\bigcup_{k\geq 1}\mathsf{MRW}_{k}(S)$, the locus of $w$ is the position on some edge $(v_{1},v_{2})$ exactly one character down from $v_{1}$ since $\mathsf{occ}_{S}(w[1..|w|-1])>\mathsf{occ}_{S}(w)$. For the sake of contradiction, we assume that the locus of $w_{1}=a_{1}u_{1}b\in\bigcup_{k\geq 1}\mathsf{MRW}_{k}(S)$ is the same as that of another $w_{2}=a_{2}u_{2}b\in\bigcup_{k\geq 1}\mathsf{MRW}_{k}(S)$ where $a_{1},a_{2},b\in\Sigma$ and $u_{1},u_{2}\in\Sigma^{\star}$. We further assume $|w_{1}|\leq|w_{2}|$ w.l.o.g. Since $a_{1}u_{1}$ and $a_{2}u_{2}$ reach the same CDAWG node, $a_{1}u_{1}$ is a suffix of $a_{2}u_{2}$. Since $w_{1}\neq w_{2}$, $|w_{1}|<|w_{2}|$ holds, and thus, $w_{1}$ is a proper suffix of $w_{2}$. This contradicts the minimality of MRW $w_{2}$. Hence, the number of MRWs in $\bigcup_{k\geq 1}\mathsf{MRW}_{k}(S)$ is never larger than $|\mathsf{E}|$. The above argument holds even if the reversal of $S$ is considered. Therefore, the upper bound is $\min\{|\mathsf{E}|,|\mathsf{E}^{R}|\}=\mathsf{e}_{\min}$. ∎
The proof of Lemma 7 implies that if the locus of an MRW $aub$ is on an edge $(v_{1},v_{2})$, then $au$ is the shortest string represented by the node $v_{1}$. Namely, $\mathsf{slink}(v_{1})=u$. From this fact, we can represent each MRW $aub$ as its corresponding edge $(v_{1},v_{2})$ where the character label of the hard Weiner link from $\mathsf{slink}(v_{1})$ to $v_{1}$ is $a$, the longest string represented by $\mathsf{slink}(v_{1})$ is $u$, and the first character of the edge label of $(v_{1},v_{2})$ is $b$. Also, by Lemma 3, we can restore the MRW $aub$ from $(v_{1},v_{2})$ in time linear in its length. Thus, the following lemma holds.
Lemma 8.
There exists a data structure of $O(\mathsf{e}_{\min})$ space that can output all MRWs occurring in $S$ in $O(|\bigcup_{k\geq 1}\mathsf{MRW}_{k}(S)|)$ time with $O(1)$ working space.
If $k$ is fixed in advance, we can store $\mathsf{MRW}_{k}(S)$ in $O(\mathsf{e}_{\min})$ space and can output them in $O(|\mathsf{MRW}_{k}(S)|)$ time. We obtain the next corollary by fixing $k=1$.
Corollary 1.
There exists a data structure of $O(\mathsf{e}_{\min})$ space that can output all MUSs for string $S$ in $O(|\mathsf{MUS}(S)|)$ time with $O(1)$ working space.
6 Conclusions and Future Work
In this paper, we proposed space-efficient $O(\mathsf{e}_{\min})$-size data structures that can output MAWs, EBFs, and MRWs for an input string $S$ of length $n$ in time linear in the output size, where $\mathsf{e}_{\min}$ denotes the minimum of the left-extensions and the right-extensions of the maximal repeats in $S$. The key tools of our method are the $O(\mathsf{e}_{\min})$-size grammar-representation of the CDAWG, and the extended longest path tree of size $O(\mathsf{e}_{\min})$.
Our future work includes the following: Is it possible to extend our method for reporting length-bounded MAWs/EBFs of Section 4.3 to the problem of reporting MAWs/EBFs of length exactly $\ell$ for a query length $\ell>1$? This would be a space-improvement of the previous work [13]. Okabe et al. [21] proposed a (non-compacted) DAWG-based data structure that can report the symmetric difference of the sets of MAWs for multiple strings, using $O(n)$ space and time linear in the output size, where $n$ is the total length of the strings. Can our CDAWG-based data structure be extended to this problem?